OpenStudy (anonymous):

Factor each polynomial: 6n^4-6

OpenStudy (anonymous):

and show work plz

OpenStudy (the_fizicx99):

is that \(\ 6n^4 - 6\) ? or \(\ 6n^{4-6}\) ?

OpenStudy (anonymous):

the first option

OpenStudy (the_fizicx99):

Yeah, find a gcf and divide it out, there's no variable gcf, but there is a coefficient gcf (6). \(\ 6(n^4 - 1) \) Now factor the differences of squares, assuming you know what they are, \(\ 6((n^2 + 1)(n - 1)(n+1)) \)

OpenStudy (anonymous):

thanks:3

OpenStudy (the_fizicx99):

Yw ;D

OpenStudy (the_fizicx99):

It seems a bit tricky in the middle, do you know how we got to n^2 +1 ?

OpenStudy (anonymous):

i think when you divide to find the gcf, you divided the exponent in half...........right?

OpenStudy (the_fizicx99):

Erm no you can't divide n^4 by 6, differences of \(\ square \), square root n^4 = n^2 and square-root of 1 is 1 (because 1* 1 = 1) Now (n^4 -1) = (n^2 + 1)(n^2 -1) = (n + 1)(n-1) 6((n^2-1)(n+1)(n-1))

OpenStudy (anonymous):

oh so like a perfect square, just for exponents?

OpenStudy (the_fizicx99):

There's a name for them, there's the perfect squares and differences of squares

OpenStudy (the_fizicx99):

They're called special products :P

OpenStudy (anonymous):

oh!

OpenStudy (the_fizicx99):

;P

OpenStudy (anonymous):

is it okay is i ask another question?

OpenStudy (the_fizicx99):

I have to get off :< it's already 9:30 and I still need to finish my assignment >.>

OpenStudy (anonymous):

oh sorry, then go, thanks for ur help and gn:3

OpenStudy (the_fizicx99):

I'll be on later if you still need help :> bai!

OpenStudy (anonymous):

baibai!!